Directions a. Please state both the null and alternative hypotheses for this question. b. Provide the decision rule for making this decision. Use an alpha level of .05. c. Show all the work necessary to calculate the appropriate statistic. d. What conclusion are you allowed to draw? Write a conclusions sentence in APA format (i.e., it includes the appropriate statistical information). A chimpanzee was trained to make same-different judgments about pairs of stimuli. For pictures of objects, the animal was correct on 23 out of 24 trials. Did the chimp perform significantly better than chance?
Answer #1

Given that,
possible chances (x)=23
sample size(n)=24
success rate ( p )= x/n = 0.958
success probability,( po )=0.5
failure probability,( qo) = 0.5
null, Ho:p=0.5
alternate, H1: p>0.5
level of significance, alpha = 0.05
from standard normal table,right tailed z alpha/2 =1.645
since our test is right-tailed
reject Ho, if zo > 1.645
we use test statistic z proportion = p-po/sqrt(poqo/n)
zo=0.95833-0.5/(sqrt(0.25)/24)
zo =4.491
| zo | =4.491
critical value
the value of |z alpha| at los 0.05% is 1.645
we got |zo| =4.491 & | z alpha | =1.645
make decision
hence value of | zo | > | z alpha| and here we reject Ho
p-value: right tail - Ha : ( p > 4.49073 ) = 0
hence value of p0.05 > 0,here we reject Ho
ANSWERS
---------------
null, Ho:p=0.5
alternate, H1: p>0.5
test statistic: 4.491
critical value: 1.645
decision: reject Ho
p-value: 0
we have enough evidence to support the claim that the chimp perform significantly better than chance